OpenStudy (anonymous):

The measures of three sides of a triangle are consecutive odd integers. The perimeter of the triangle is 57. What are the measures of the three sides?

OpenStudy (anonymous):

Let teh numbers be A,B,C then we can assign values A=x B=x+2 C=x+4 we know A+B+C=57 by definition of perimiter, so then x+ (x+2) + (x+4) = 57 3x = 57-6 x = 51/3 x= 17 now then A =17, B=19, C=21